Fermi National Laboratory
 

Accelerator Update for Jun 8, 2006 - June 9, 2006

 

Thursday June 8

The day shift began with Operations monitoring the Recycler stash, with the Antiproton Source (Pbar) not stacking due to problems with an Accumulator Septa (A:ISEP2V), with the Booster, Main Injector (MI), Recycler, and Switchyard (SY), Tevatron (TeV) conducting studies, and with NuMI and MiniBooNE taking beam.

Operations prepared Pbar for access at 9:15 AM to continue with a ground fault investigation for an Accumulator magnet (A:QT).

Operations prepared D-Zero (D0) for access at 10:17 AM. The Collider Detector at Fermilab (CDF) will use this as quiet time for their experiment.

Operations reported at 10:20 AM that a vacuum leak in the Tevatron Electron Lens (TEL2) was getting worse. Vacuum techs accessed the TeV, added a vacuum pump and oil to another pump and then came out 11:51 AM.

At 1:02 PM, MI suffered a vacuum burst in the MI-8 line. Operations let the vacuum recover and then turned MI back about ten minutes later.

Operations put the TeV into a dry squeeze at 1:24 PM.

A Septa expert reported at 1:58 PM that the problem with A:ISEP was found.

Operations established beam in the TeV at 2:08 PM. Experts began a study period.

At 3:25 PM, Booster experts reported that work was beginning on the Klystron Debuncher water skid. This held off all downstream beam until 3:52 PM.

At 3:56 PM, the TeV aborted. There was no quench. Experts investigated and found that TECAR had pulled the abort. All the VCBs opened. Investigations continue.

Work on A:QT finished and Pbar attempted to stack at 6:34 PM.

At 7:28 PM, A:QT tripped off. Operations called in an expert.

At 8:25 PM, the Klystron Debuncher cavity water skid tripped off for the first of three times in fifty minutes. It reset each time.

Operations reported at 9:20 PM that A:QT was operating normally. Pbar began stacking at 10:07 PM.

Friday June 9

The midnight shift began with Operations monitoring the stash, with Pbar stacking, with the TeV, MI, RR, and SY conducting studies, and with MiniBooNE and NuMI taking beam.

At 2:21 AM, Operations reported that the Klystron Debuncher water skid tripped off for the second of five times during the midnight shift. It reset each time.

At 4:55 AM, the TeV quenched at sector F1. Operations discovered that beam was injected at the wrong time.

Operations put the TeV into a dry squeeze at 6 AM. Studies began at 6:26 AM.

The Plans for Friday and the Weekend

The plan for today is to establish a small store in the TeV this afternoon. The weekend will consist of commissioning.

Machine Reports

Linac

Experts have added diagnostics to the Klystron Debuncher cavity water skid.

Booster

BRF1 still has problems. Experts hope to bring BRF19 back on line today.

Main Injector

Experts report that the machine is running smooth.

Antiproton Source

Experts reported that they’re stacking at 4-5mA/hour.

TeV

Experts believe that the ramp dump was due to a current monitor (DCCT) and a sticky button. They are ready for antiprotons.

Recycler

Experts are ready to help establish small TeV store.

Switchyard 120

Experts are commissioning power supplies

MiniBooNE

Experts are trying to resolve some problems with their horn.

NuMI

Experts have been trying to interleave stacking and non-stacking time events.

Collider Detector at Fermilab

Experimenters used the D0 access yesterday as quiet time. They have a problem with their detector that needs an access to repair.

D-Zero

Experimenters took cosmic rays overnight.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/09/2006    email Fermilab